# STAR-RIS Enhanced UAV-Enabled MEC Networks with Bi-Directional Task Offloading

### v1

@@ III-B3 Designing $\boldsymbol{\Xi}_{3}$ with given $\boldsymbol{\zeta}$ and $\boldsymbol{\Upsilon}$
III-C Proposed Optimization Algorithm & Analysis on Complexity and Convergence
The presented iterative algorithm for solving the original optimization problem (III-A) is summarized as Algorithm 2, which is a block coordinate descent (BCD) algorithm designed to tackle the three subproblems explained in Section III. The value $v$ represents the difference of the obtained objective function value between adjacent iterations. Once $v$ falls below a predefined threshold $\varepsilon$, the proposed algorithm will converge.
Algorithm 2: Proposed Iterative Algorithm to Handle the Optimization Problem (III-A)
@@ III-C Proposed Optimization Algorithm & Analysis on Complexity and Convergence
   index $\widehat{l}$ = 0.
2: While $v>\varepsilon$ or $\widehat{l}=0$ do
3:  Solve the sub-problem (III-B1) by utilizing Algorithm 1
  with the given $\boldsymbol{\Upsilon}^{(\widehat{l})}$ and $\mathbf{Q}^{(\widehat{l})}$, and update $\mathbf{L}^{(\widehat{l}+1)}$ and
  $\boldsymbol{\zeta}^{(\widehat{l}+1)}$ with the obtained solution.
4:  Solve the sub-problem (III-B2) based on the Dinkelbach’s
  algorithm with the given $\boldsymbol{\zeta}^{(\widehat{l}+1)}$ and $\mathbf{Q}^{(\widehat{l})}$, and update
  $\mathbf{L}^{(\widehat{l}+1)}$ and $\boldsymbol{\Upsilon}^{(\widehat{l}+1)}$ with the obtained solution.
5:  Solve the sub-problem (III-B3) based on the Dinkelbach’s
  algorithm with the given $\boldsymbol{\zeta}^{(\widehat{l}+1)}$ and $\boldsymbol{\Upsilon}^{(\widehat{l}+1)}$, and update
  $\mathbf{L}^{(\widehat{l}+1)}$ and $\mathbf{Q}^{(\widehat{l}+1)}$ with the obtained solution.
6: Calculate the objective value $\overline{R}^{(\widehat{l}+1)}=\frac{L_{\mathrm{tol}}^{(\widehat{l}+1)}}{E_{\mathrm{tol}}^{(\widehat{l}+1)}}$ and
  update $v=$$\left|\overline{R}^{(\widehat{l}+1)}-\overline{R}^{(\widehat{l})}\right|$ based the obtained solu-
  tions; Let $\widehat{l}=\widehat{l}+1$.
@@
Then, the computational complexity of the proposed iterative algorithm is analysed. Specifically, the computational complexity mainly comes from solving the divided three subproblems, i.e., (III-B1), (III-B2) and (III-B3). Regarding the first subproblem, we propose a two-tier algorithm that incorporates both Dinkelbach’s algorithm and the penalty method to effectively address it. It is presumed that the interior point method will be employed to compute the transformed standard convex optimization problem. Consequently, the computational complexity for resolving the first subproblem can be expressed as $\mathcal{O}_{1}=\mathcal{O}(L_{1}L_{2}(5NK)^{3.5})$, where $L_{1}$ and $L_{2}$ denote the number of iterations in the outer and inner loops, respectively. In terms of the subproblem designing the passive beamforming, the computational complexity is dominated by handling the optimization problem (III-B2), which is given by $\mathcal{O}_{2}=\mathcal{O}(L_{3}(4NK+2M)^{3.5})$, where $L_{3}$ is the number of iteration. For the subproblem 3, the SCA technique is adopted to transform the UAV trajectory problem into a convex optimization problem. Supposing $L_{4}$ is the number of iteration, the the computational complexity can be determined as $\mathcal{O}_{3}=\mathcal{O}(L_{4}(4NK+2N)^{3.5})$. Therefore, the total computing complexity of the proposed algorithm can be expressed as $\mathcal{O}_{\mathrm{tol}}=L(\mathcal{O}_{1}+\mathcal{O}_{2}+\mathcal{O}_{2})$, where $L$ denotes the total iteration number of the proposed algorithm. Based on the overall computational complexity analysis, it is evident that the computing complexity is intricately connected to both the quantity of sub-time slots ($N$) and the number of elements installed at STAR-RIS ($M$).
Actually, the convergence of the proposed algorithm can be guaranteed by utilizing the BCD algorithm to address the optimization problem (III-A). Because this approach ensures that each iteration results in a solution that is at least as good as the previous one and thus the objective function is monotonically non-decreasing versus the iteration. Additionally, we will also verify the convergence of the proposed algorithm through the simulation results presented in Section IV.
@@ IV Simulation Results
To highlight the effectiveness of the proposed STAR-RIS-aided UAV-enabled MEC scheme, we present the numerical simulation results in this section and compare the their with three benchmark schemes, including: 1) RIS-aided scheme [23]: In this baseline scheme, two adjacent conventional RISs with $\frac{M}{2}$ elements are adopted to replace the STAR-RIS, where one is the reflecting-only RIS and the other one is the transmission-only RIS; 2) Without trajectory optimization scheme: This scheme focuses on optimizing $\mathbf{L}$, $\boldsymbol{\zeta}$ and $\boldsymbol{\Upsilon}$ with direct UAV trajectory flying from the initial point to the final point at a consistent speed. 3) Heuristic scheme: UAV will traverse each user node based on the pre-defined trajectory at a consistent speed. Similarly, variables $\mathbf{L}$, $\boldsymbol{\zeta}$ and $\boldsymbol{\Upsilon}$ will be optimized in this scheme. In addition, the simulation parameters are listed in Table I.
@@
Altitude of UAV
$H=30$ m
Bandwidth
$B=1$ MHz
Carrier frequency
$\lambda_{\mathrm{c}}=2.4$ GHz
Effective capacitance coefficient of MEC servers
$\iota_{\mathrm{ua}}=\iota_{\mathrm{BS}}=10^{-27}$
Initial/final point of UAV trajectory
$\mathbf{q}_{\mathrm{I}}=[-40,0,30]^{T}$ m, $\mathbf{q}_{\mathrm{F}}=[40,0,30]^{T}$ m
Maximum flight velocity and acceleration of UAV
[…]
Maximum CPU frequency
$F_{\mathrm{BS}}=20$ GHz, $F_{\mathrm{ua}}=12$ GHz
Noise power
$\sigma^{2}_{\mathrm{ua}}=\sigma^{2}_{\mathrm{BS}}=-100$ dBm
Pass loss exponent
$\alpha_{\mathrm{r}k}=2.3$, $\alpha_{\mathrm{rb}}=2.3$
Scaling factor
$\omega=10$
The number of CPU cycles needed for processing 1-bit of data
$\varrho_{\mathrm{ua}}=\varrho_{\mathrm{BS}}=10^{3}$ cycles/bit
Time slot
$\delta_{\mathrm{t}}=0.2$ s
Transmitting power at users
$p=20$ dBm,
Tolerance accuracy thresholds
$\varepsilon=\varepsilon_{1}=\varepsilon_{2}=10^{-3}$
Figure 2: Energy efficiency versus iterations index with $L_{k}=20$ Mbits, with different mission period $T$ and the number of elements at STAR-RIS $M$.
@@
In Fig. 3 and Fig. 4, we give the UAV trajectories considering different parameter settings, where $M$ is fixed as 36. Specifically, UAV trajectories under different mission period $T$ and time slot number $N$ are shown in Fig. 3, where all the users with computation requirements $L_{k}=20$Mbits. We can find that the UAV consistently flies towards User 3, who is the furthest user away from the BS, in order to improve the channel quality between User 3 and the BS under different $T$. This is done to ensure that User 3’s minimum task requirement is met. Additionally, it demonstrates that the UAV tends to approach the BS as $T$ increases, which aims to offload larger tasks to the BS, ultimately enhancing the energy efficiency. In Fig. 4, the UAV trajectories is plotted with different $\{L_{k}\}_{k\in\mathcal{K}}$. It is observed that when all users share the same minimum task requirement, the UAV prioritizes traversing users located at longer distances from the BS, such as User 1, User 2, and User 3. However, if certain users, like User 3 and User 5, require a higher number of offloading task bits, the UAV will approach these users to improve their channel quality and fulfil their task requirements.
@@
Figure 3: UAV trajectory with $M=36$ and $L_{k}=20$ Mbits, as well as the different mission period $T$ and $N$.
Figure 4: UAV trajectory with different $L_{k}$, with $T=10$ s, $N=50$ and $M=36$.
@@
Figure 5: Energy efficiency versus the number of elements at STAR-RIS with $N=50$ and $L_{k}=20$ Mbits, $\forall k\in\mathcal{K}$.
@@
Next, we investigate the influence of the number of elements equipped at STAR-RIS, i.e., $M$, on the energy efficiency with $N=50$ and $L_{k}=20$ Mbits, $\forall k\in\mathcal{K}$. Specifically, it can be observed that the energy efficiency of all the schemes increases as $M$ grows, as the additional elements can offer more flexibility to reconfigure the wireless environment. However, the rates of increase gradually decrease as $M$ continues to grow. The proposed scheme offers a greater performance gain in improving the energy efficiency, especially when the number of elements is limited, compared to the conventional RIS-assisted scheme. Additionally, an interesting observation is obtained between the RIS-aided scheme and the heuristic scheme. Specifically, the heuristic scheme achieves higher energy efficiency than the RIS-aided scheme when the $M$ is small. As $M$ increases, the RIS-aided scheme becomes more energy efficient than the heuristic scheme. This indicates that the trajectory of the UAV, with the help of the RIS, has the potential to overcome performance limitations imposed by other system settings. These observations are also evident when comparing with the heuristic scheme.
@@
Figure 6: Energy efficiency versus the mission period $T$ with $M=50$ and $L_{k}=20$ Mbits, $\forall k\in\mathcal{K}$.
@@
Fig. 6 explores the impact of the mission duration $T$ on energy efficiency when $M$ is fixed at $50$ and $L_{k}$ is set to $20$ Mbits $\forall k\in\mathcal{K}$. The results show that the proposed scheme provides the highest performance gain for MEC network, while the scheme without the trajectory optimization exhibits the lowest energy efficiency. Regarding the proposed scheme, the energy efficiency consistently increases as $T$ grows from $8$s to $18$s. This can be attributed to the fact that the UAV having more time to optimize its trajectory, thereby improving the channel condition between users/BS and UAV as $T$ increases. However, beyond $18$s, the energy efficiency starts to decline. This is due to the channel quality between users/BS and UAV reaching a saturation point, which may not increase the completed task bits but increase the energy consumption of UAV as $T$ becomes larger. Besides, the conventional RIS-assisted scheme has the similar trend in energy efficiency as the proposed scheme. However, the energy efficiency reaches the limitation at $T=12$s in this scheme, which indicates that STAR-RIS demonstrates a significant potential in achieving a balance between system energy consumption and the volume of the offloaded data when compared to the conventional RIS. In contrast to the suggested scheme, both the heuristic scheme and the scheme with direct UAV trajectory suffer from the earlier performance limitation as $T$ increases. This highlights the significance of the UAV’s trajectory optimization in augmenting the energy efficiency of the comprehensive UAV-enabled MEC system.

### v3

@@ III-B3 Designing $\boldsymbol{\Xi}_{3}$ with given $\boldsymbol{\zeta}$ and $\boldsymbol{\Upsilon}$
III-C Proposed Optimization Algorithm Analysis
The presented iterative algorithm for solving the original optimization problem (26) is summarized as Algorithm 2, which is two-tier iterative algorithm designed to tackle the three subproblems explained in Section III. Specifically, the inner loop is responsible for solving the binary variable, i.e., user association variable, by progressively increase the penalty coefficient, $\widehat{\rho}$. $v$ represents the objective function value obtained after the transformation through the Dinkelbach’s algorithm. Once $v$ falls below a predefined threshold $\varepsilon$, the proposed algorithm will converge.
Then, the computational complexity of the proposed iterative algorithm is analysed. Specifically, the computational complexity mainly comes from solving the divided three subproblems, i.e., (29), (34) and (39). Regarding the first subproblem, we propose an iterative penalty algorithm that incorporates both Dinkelbach’s algorithm to effectively address it. It is presumed that the interior point method will be employed to compute the transformed standard convex optimization problem. Consequently, the computational complexity for resolving the first subproblem can be expressed as $\mathcal{O}_{1}=\mathcal{O}(L_{1}(5NK)^{3.5})$, where $L_{1}$ denotes the total iterations number of the Algorithm 1. The computational complexity of addressing the subproblem (38) can be expressed as $\mathcal{O}_{2}=\mathcal{O}((6NK+2M)^{3.5})$, which is significantly lower than the computing complexity associated with using the semidefinite relaxation (SDR) method for addressing this subproblem, given by $\mathcal{O}((4NK+2M^{2})^{3.5})$. For the subproblem 3, the SCA technique is adopted to transform the UAV trajectory problem into a convex optimization problem. The the computational complexity can be determined as $\mathcal{O}_{3}=\mathcal{O}((4NK+2N)^{3.5})$. Therefore, the total computing complexity of the proposed algorithm can be expressed as $\mathcal{O}_{\mathrm{tol}}=\widehat{L}(\mathcal{O}_{1}+\mathcal{O}_{2}+\mathcal{O}_{2})$, where $\widehat{L}$ denotes the total iteration number of the proposed algorithm. Based on the overall computational complexity analysis, it is evident that the computing complexity is intricately connected to both the quantity of sub-time slots ($N$) and the number of elements installed at STAR-RIS ($M$).
Algorithm 2: Proposed Iterative Algorithm to Handle the Optimization Problem (26)
@@ III-C Proposed Optimization Algorithm Analysis
   index $l$ = 0; Calculate $\psi^{(0)}$ with the given initial feasible
   point.
2: While $v>\varepsilon$ or ł$\l=0$ do
3:  Solve the sub-problem (29) by utilizing Algorithm 1
  with the given $\boldsymbol{\Upsilon}^{(l)}$ and $\mathbf{Q}^{(l)}$, and update $\mathbf{L}^{(l+1)}$ and
  $\boldsymbol{\zeta}^{(l+1)}$ with the obtained solution.
4:  Solve the sub-problem (38) with the given $\boldsymbol{\zeta}^{(l+1)}$ and
  $\mathbf{Q}^{(l)}$, and update $\mathbf{L}^{(l+1)}$ and $\boldsymbol{\Upsilon}^{(l+1)}$ with the obtained
  solution.
5:  Solve the sub-problem (48) with the given $\boldsymbol{\zeta}^{(l+1)}$ and
  $\boldsymbol{\Upsilon}^{(l+1)}$, and update $\mathbf{L}^{(l+1)}$ and $\mathbf{Q}^{(l+1)}$ with the obtai-
  ned solution.
6:  Calculate $v=L_{\mathrm{tol}}^{(l+1)}-\psi^{(l)}E_{\mathrm{tol}}^{(l+1)}$ based on the obtain-
  ed solutions; Update $\psi^{(l+1)}=\frac{L_{\mathrm{tol}}^{(l+1)}}{E_{\mathrm{tol}}^{(l+1)}}$; Let $l=l+1$.
@@
Actually, the convergence of the proposed algorithm can be guaranteed, since the framework of the block coordinate descent (BCD) algorithm is utilized to address the optimization problem (26), which will ensure that each iteration results in a solution that is at least as good as the previous one and thus the objective function is monotonically non-decreasing versus the iteration. Additionally, we will also verify the convergence of the proposed algorithm through the simulation results presented in Section IV. To evaluate the quality of the solutions obtained from the proposed algorithm, the SDR method is selected as the algorithm for comparison in Section IV, which is commonly utilized for addressing optimization challenges in the field of communication. Since the approximation accuracy of solutions obtained through the SDR method in solving various types of optimization problems has been theoretically demonstrated [46]. In addition, the authors of [46] thoroughly summarize the theoretical approximation accuracy of solving various optimization problems in the field of communication using SDR methodology. Note that the approximation accuracy represents the ratio between the obtained solution by leveraging the SDR and the theoretically optimal solution. Thus, choosing the SDR method as a comparative benchmark algorithm can provide an effective evaluation of the theoretical gap between the achieved solution and the optimal solution.
@@ IV Simulation Results
To highlight the effectiveness of the proposed STAR-RIS-aided UAV-enabled MEC scheme, we present the numerical simulation results in this section and compare the their with three benchmark schemes, including: 1) RIS-aided scheme [33]: In this baseline scheme, two adjacent conventional RISs with $\frac{M}{2}$ elements are adopted to replace the STAR-RIS, where one is the reflecting-only RIS and the other one is the transmission-only RIS; 2) Fixing trajectory scheme: This scheme focuses on optimizing $\mathbf{L}$, $\boldsymbol{\zeta}$ and $\boldsymbol{\Upsilon}$ with direct UAV trajectory flying from the initial point to the final point at a consistent speed. 3) Heuristic scheme: UAV will traverse each user node based on the pre-defined trajectory at a consistent speed. Similarly, variables $\mathbf{L}$, $\boldsymbol{\zeta}$ and $\boldsymbol{\Upsilon}$ will be optimized in this scheme. 4) SDR scheme: The SDR method is utilized to optimize the passive beamforming variable $\boldsymbol{\Upsilon}$ in this scheme. In addition, the simulation parameters are listed in Table I.
@@
Altitude of UAV $H=30$ m
Bandwidth $B=1$ MHz
Carrier frequency $\lambda_{\mathrm{c}}=2.4$ GHz
Effective capacitance coefficient of MEC servers $\iota_{\mathrm{ua}}=\iota_{\mathrm{BS}}=10^{-27}$
Initial/final point of UAV trajectory $\mathbf{q}_{\mathrm{I}}=[-40,0,30]^{T}$ m, $\mathbf{q}_{\mathrm{F}}=[40,0,30]^{T}$ m
Maximum flight velocity and acceleration of UAV $v_{\max}=30$ m/s, $a_{\max}=20$ m/s2
Maximum CPU frequency $F_{\mathrm{BS}}=20$ GHz
Noise power $\sigma^{2}_{\mathrm{ua}}=\sigma^{2}_{\mathrm{BS}}=-100$ dBm
Pass loss exponent $\alpha_{\mathrm{r}k}=2.3$, $\alpha_{\mathrm{rb}}=2.3$
Scaling factor $\omega=10$
The number of CPU cycles needed for processing 1-bit of data $\varrho_{\mathrm{ua}}=\varrho_{\mathrm{BS}}=10^{3}$ cycles/bit
Time slot $\delta_{\mathrm{t}}=0.2$ s
Transmitting power at users $p=20$ dBm
Tolerance accuracy thresholds $\varepsilon=\varepsilon_{1}=\varepsilon_{2}=10^{-3}$
Figure 3: Energy efficiency versus iterations index with $L_{k}=20$ Mbits and $F_{\mathrm{ua}}=12$ GHz, with different mission period $T$ and the number of elements at STAR-RIS $M$.
@@
Figure 4: UAV trajectory with $M=36$, $L_{k}=20$ Mbits, $\forall k\in\mathcal{K}$ and $F_{\mathrm{ua}}=12$ GHz, as well as the different mission period $T$ and $N$.
In Fig. 4 and Fig. 5, we give the UAV trajectories considering different parameter settings, where $M$ and $F_{\mathrm{ua}}$ are fixed as $36$ and $12$ GHz, respectively. Specifically, UAV trajectories under different mission period $T$ and time slot number $N$ are shown in Fig. 4, where all the users with computation requirements $L_{k}=20$Mbits. We can find that the UAV consistently flies towards User 3, who is the furthest user away from the BS, in order to improve the channel quality between User 3 and the BS under different $T$. This is done to ensure that User 3’s minimum task requirement is met. Additionally, it demonstrates that the UAV tends to approach the BS as $T$ increases, which aims to offload larger tasks to the BS, ultimately enhancing the energy efficiency. In Fig. 5, the UAV trajectories is plotted with different $\{L_{k}\}_{k\in\mathcal{K}}$. It is observed that when all users share the same minimum task requirement, the UAV prioritizes traversing users located at longer distances from the BS, such as User 1, User 2, and User 3. However, if certain users, like User 3 and User 5, require a higher number of offloading task bits, the UAV will approach these users to improve their channel quality and fulfil their task requirements.
Figure 5: UAV trajectory with different $L_{k}$, with $T=10$ s, $N=50$, $F_{\mathrm{ua}}=12$ GHz and $M=36$.
@@
Figure 6: Energy efficiency versus the number of elements at STAR-RIS with $N=50$, $F_{\mathrm{ua}}=12$ GHz and $L_{k}=20$ Mbits, $\forall k\in\mathcal{K}$.
@@
Next, we investigate the influence of the number of elements equipped at STAR-RIS, i.e., $M$, on the energy efficiency with $N=50$, $F_{\mathrm{ua}}=12$ GHz and $L_{k}=20$ Mbits, $\forall k\in\mathcal{K}$, as shown in Fig. 6. Specifically, it can be observed that the energy efficiency of all the schemes increases as $M$ grows, as the additional elements can offer more flexibility to reconfigure the wireless environment. However, the rates of increase gradually decrease as $M$ continues to grow. The proposed scheme offers a greater performance gain in improving the energy efficiency, especially when the number of elements is limited, compared to the conventional RIS-assisted scheme. Additionally, an interesting observation is obtained between the RIS-aided scheme and the heuristic scheme. Specifically, the heuristic scheme achieves higher energy efficiency than the RIS-aided scheme when the $M$ is small. As $M$ increases, the RIS-aided scheme becomes more energy efficient than the heuristic scheme. This indicates that the trajectory of the UAV, with the help of the RIS, has the potential to overcome performance limitations imposed by other system settings. The presented results in fixing trajectory scheme further verify the importance of the UAV’s trajectory in enhancing the system performance. Note that, the proposed scheme illustrates slight advantages through a comparison with the results obtained by the SDR scheme, showcasing the effectiveness of the proposed algorithm.
@@
Figure 7: Energy efficiency versus the mission period $T$ with $M=50$, $F_{\mathrm{ua}}=12$ GHz and $L_{k}=20$ Mbits, $\forall k\in\mathcal{K}$.
@@
Fig. 7 explores the impact of the mission duration $T$ on energy efficiency when $M$, $F_{\mathrm{ua}}$ and $L_{k},\forall k\in\mathcal{K}$ are respectively set to $50$, $12$ GHz and $20$ Mbits. The results show that the proposed scheme provides the highest performance gain for MEC network, while the scheme without the trajectory optimization exhibits the lowest energy efficiency. Regarding the proposed scheme and the SDR scheme, the energy efficiency consistently increases as $T$ grows from $8$s to $16$s. This can be attributed to the fact that the UAV having more time to optimize its trajectory, thereby improving the channel condition between users/BS and UAV as $T$ increases. However, beyond $16$s, the energy efficiency starts to decline. This is due to the channel quality between users/BS and UAV reaching a saturation point, which may not increase the completed task bits but increase the energy consumption of UAV as $T$ becomes larger. Besides, the conventional RIS-assisted scheme has the similar trend in energy efficiency as the proposed scheme. However, the energy efficiency reaches the limitation at $T=12$s in this scheme, which indicates that STAR-RIS demonstrates a significant potential in achieving a balance between system energy consumption and the volume of the offloaded data when compared to the conventional RIS. In contrast to the suggested scheme, both the heuristic scheme and the scheme with direct UAV trajectory suffer from the earlier performance limitation as $T$ increases. This highlights the significance of the UAV’s trajectory optimization in augmenting the energy efficiency of the comprehensive UAV-enabled MEC system. Furthermore, we can find that the outcomes attained by the suggested algorithm consistently surpass those of the SDR scheme, underscoring the potential and benefits of the proposed algorithm.
Figure 8: Energy efficiency versus the CPU frequency of the MEC server installed at the UAV, $F_{\mathrm{ua}}$, with $M=50$, $N=50$ and $L_{k}=10$ Mbits, $\forall k\in\mathcal{K}$.
Figure 8 illustrates the fluctuating energy efficiency trend corresponding to the CPU frequency of the MEC server deployed on the UAV with parameters $M=50$, $N=50$, and $L_{k}=10$ Mbits, $\forall k\in\mathcal{K}$. In particular, the escalation of $F_{\mathrm{ua}}$ correlates with a progressive enhancement in energy efficiency across all scenarios, characterized by a diminishing rate of increase. There are two reasons for this: (i) By elevating the CPU frequency of the MEC server located at the UAV, users can delegate an expanded array of computational duties to UAV servers for execution. Although this heightened task allocation could potentially elevate energy consumption levels, the pace at which tasks are added outpaces the growth in energy usage, consequently bolstering the system’s overall energy efficiency. (ii) The limitation of achievable offloading rate for users is influenced by system parameters such as allocated task time, number of RIS elements, and users’ transmit power. Additionally, the energy consumption of computing tasks assigned to UAVs will increase in a cubic function manner as more tasks are delegated (see (23)), contributing to a gradual slowdown in energy efficiency improvement. Furthermore, when compared to the four MEC schemes incorporating trajectory design, i.e., the proposed scheme, RIS-aided scheme, Heuristic scheme and SDR scheme, there is a gradually increase performance gap observed with the scheme lacking UAV trajectory optimization with the growth of $F_{\mathrm{ua}}$. This indicates that the trajectory optimization plays an important role in balancing the computation bits and energy consumption.
Figure 9: Energy efficiency versus CPU cycles required at the UAV for computing 1-bit of task-input data with $M=50$, $N=50$, $F_{\mathrm{ua}}=12$ GHz and $L_{k}=20$ Mbits, $\forall k\in\mathcal{K}$.
The impact of the CPU processing cycles needed by the UAV to compute a single bit of task input data, denoted as $\varrho_{\mathrm{ua}}$, on energy efficiency is investigated with considerations for $M=50$, $N=50$, and $L_{k}=20$ Mbits, $\forall k\in\mathcal{K}$. Specifically, it is observed that the energy efficiency declines progressively as the parameter denoted by $\varrho_{\mathrm{ua}}$ increases for all scenarios. This is due to the fact that the MEC server installed on the UAV must assign additional frequency resources to handle 1-bit data, leading to increased energy consumption for executing offloaded tasks from users. Furthermore, the MEC scheme lacking the optimization of the UAV trajectory demonstrates the least energy efficiency compared with the other four schemes, which highlights the significant potential of optimizing the UAV trajectory design to enhance the energy efficiency of the MEC system. It is worth noting that the SDR scheme exhibits the highest performance gain by comparing the obtained simulation results with other four schemes when the $\varrho_{\mathrm{ua}}\leq 950$. However, the benefit is exceedingly marginal when contrasted with the proposed scheme. The energy efficiency of the SDR scheme diminishes rapidly as the value of $\varrho_{\mathrm{ua}}$ increases beyond $950$, with the outcomes achieved even falling below those of the Heuristic scheme when $\varrho_{\mathrm{ua}}$ ranges from $1150$ to $1300$. The results of the simulation demonstrate a consistent and stable reduction when employing the suggested approach, surpassing the performance of the RIS-aided scheme, the fixed trajectory scheme, and the heuristic scheme. This indicates the high robustness of the proposed algorithm.
(a) The allocation of the offloaded bits at the UAV and the BS versus the time slot.
(b) User scheduling for task offloading versus the time slot.
Figure 10: The allocation of the offloaded tasks and user scheduling with $M=50$, $F_{\mathrm{ua}}=6$ GHz, $N=50$ and $L_{k}=15$ Mbits, $\forall k\in\mathcal{K}$.
Finally, to reveal the collaboration between the UAV and the BS for simultaneously processing the offloaded tasks from users, Fig. 10(10(a)) details the distribution of offloaded bits at the UAV and the BS throughout the mission duration $T$. Specifically, it is noted that the MEC server mounted on the UAV is responsible for processing the majority of computing tasks initiated by users. To meet the demanding computational needs of users, the UAV must primarily handle the processing of offloaded tasks, which is because the UAV’s mobility and flexibility are beneficial for enhancing the channel quality between users and the UAV, and the two-path loss of the offloading signals restricts users’ ability to offload tasks to the BS. Additionally, we can find that the dynamic trend is evident in the distribution of the offloaded bits at the UAV and the BS, which is corresponding to the UAV’s mobility. To show more details, the user scheduling in mission period is presented in Fig. 10(10(b)). In particular, it is observed that the considered five users (from User 1 to User 5) are allocated $10$, $6$, $6$, $15$, $12$ time slots to offload their computing tasks to the MEC servers situated at the UAV and the BS, respectively. Note that the final time slot remains unassigned for task offloading, in accordance with the anticipated scheduling. By comparing Fig. 10(10(a)) and 10(10(b)), it is important to highlight that the individuals allocated with more time slots, such as User 4 and User 5, are inclined to reduce the computational bits transferred to the UAV while opting to increase the computational bits directed to the BS within each time slot, which is beneficial for increasing the energy efficiency of the MEC system. To fulfil the minimal task offloading requirement, certain users such as User 2 and User 3, who possess limited time slots, opt to augment the volume of offloaded tasks directed to the MEC server situated on the UAV.